OpenStudy (anonymous):

what is 5 2/3 divided by 1 5/6

OpenStudy (anonymous):

First convert the two mixed fractions into improper fractions. 5 2/3 = 5 + 2/3 = ? 1 5/6 = 1 + 5/6 = ? Then, you can divide the resulting fractions,